a body of mass 1 kg is suspended from the free end of a spring having force constant 400n/m is executing s.h.m.when the total energy of the system is 2 joule,find the maximum acceleration

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
w = sqrt(k/m) = 20
1/2 m w² A² = 1/10
 
 
 
Acceleration = A w² = 400/10 = 40
 
Hopw it helps
 
Thanks and regards
Vikas TU
4 years ago
Dear student 
For more such examples nd concept please refer the link below 
Good Luck